A die is thrown twice and the sum of the numbers rising is noted to be 6. Calculate the is the conditional probability that the number 4 has arrived at least once?

Grade:12

1 Answers

SJ
3 years ago
Welcome to Askiitians

If a dice is thrown twice, then the sample space obtained is:

S = {(1,1)(1,2)(1,3)(1,4)(1,5)(1,6)

(2,1)(2,2)(2,3)(2,4)(2,5)(2,6)

(3,1)(3,2)(3,3)(3,4)(3,5)(3,6)

(4,1)(4,2)(4,3)(4,4)(4,5)(4,6)

(5,1)(5,2)(5,3)(5,4)(5,5)(5,6)

(6,1)(6,2)(6,3)(6,4)(6,5)(6,6)}

From the given data, it is needed to find the Probability that 4 has appeared at least once, given the sum of nos. is observed to be 6

Assume that, F: Addition of numbers is 6

and take E: 4 has appeared at least once

So, that, we need to find P(E|F)

Finding P (E):

The probability of getting 4 atleast once is:

E = {(1, 4),(2, 4),(3, 4),(4, 4),(5, 4),(6, 4),(4, 1),(4, 2),(4, 3),(4, 5),(4, 6)}

Thus , P(E) = 11/ 36

Finding P (F):

The probability to get the addition of numbers is 6 is:

F = {(1, 5), (5, 1), (2, 4), (4, 2), (3, 3)}

Thus, P(F) = 5/ 36

Also, E∩F = {(2,4), (4,2)}

P(E∩F) = 2/36

Thus, P(E|F) = (P(E∩F) ) / (P (F) )

Now, subsbtitute the probability values obtained= (2/36)/ (5/36)

Hence, Required probability is 2/5.

Thanks